### v2

# A Real-time Dyson Expansion Scheme: Efficient Inclusion of Dynamical Correlations in Non-equilibrium Spectral Properties

Abstract
Time-resolved photoemission spectroscopy is the key technique to probe the real-time non-equilibrium dynamics of electronic states. Theoretical predictions of the time dependent spectral function for realistic systems is however, a challenge. Employing the Kadanoff-Baym equations to find this quantity results in a cubic scaling in the total number of time steps, quickly becoming prohibitive and often fail quantitatively and even qualitatively. In comparison, mean-field methods have more favorable numerical scaling both in the number of time steps and in the complexity associated with the cost of evolving for a single time step, however they miss key spectral properties such as emergent spectral features. Here we present a scheme that allows for the inclusion of dynamical correlations to the spectral function while maintaining the same scaling in the number of time steps as for mean-field approaches, while capturing the emergent physics. Further, the scheme can be efficiently implemented on top of equilibrium real-time many-body perturbation theory schemes and codes. We see excellent agreement with exact results for test systems. Furthermore we exemplify the method on a periodic system and demonstrate clear evidence that our proposed scheme produces complex spectral features including excitonic band replicas, features that are not observed using static mean-field approaches.
preprint: APS/123-QED
Introduction
In recent years there has been great interest in the non-equilibrium properties of quantum materials. These systems can display phenomena such as ultrafast phase transitions[1, 2], formation of exotic states[3, 4] and allow tuning of material properties through driving[5, 6]. The first principles description of non-equilibrium dynamics of electrons, motivated by advances in experiments [7, 8, 9], has attracted attention in recent years [10, 11, 12] and stimulated new theoretical developments [13, 14, 15]. In particular the Hartree-Fock Generalized Kadanoff-Baym ansatz (HF-GKBA) emerged as a powerful tool[16, 13, 10]. It has been shown in many scenarios to provide results that are as accurate as the Kadanoff-Baym equations (KBEs)[17, 18, 19, 20, 21], which HF-GKBA approximates. Unlike the KBEs, the HF-GKBA does not require computation of the Green’s function (GF) on the full $t,t^{\prime}$ grid (Fig 1)[13, 22]. Further, KBEs can suffer from spurious damping[23, 24]. There is however one major caveat: these excellent results are restricted to properties derived from the time-diagonal component of the Green’s function. Since the HF-GKBA reconstructs the time off-diagonal components at the Hartree-Fock (HF) level, properties linked to these components perform similarly to HF results[25]. In some cases the HF-GKBA results are even worse than HF alone[26].
Time off-diagonal components are however, the key ingredient in computing the non-equilibrium time-resolved spectral function[27]. This gives information about the non-equilibrium energy distribution of quasiparticles in the system, and captures exciton formation and non-equilibrium phase transitions[28, 29, 2]. Furthermore, to see the intra-gap electron-hole signatures it is necessary to include two-body correlations when computing the time off-diagonals. A feature which we will show is missed by HF and the HF-GKBA.
Figure 1: Schematic of the off-diagonal GKBA scheme. In the two step procedure one first performs a time evolution along the time-diagonal. The information from the diagonal is then used as a starting point to update the off-diagonal
While methods have been employed to improve the HF-GKBA, they have mainly been applied to improve the time-diagonal and steady-state properties[30, 31]. Works that seek to improve the spectral properties[32] are scarce and the improvements do not go beyond the mean-field level due to the underlying reconstruction of the off-diagonal. Performing the full KBE calculations[33, 34] instead is however problematic as the computational scaling is cubic in the number of time-steps, making it numerically impractical. Although methods exist to reduce the scaling of the KBEs[14, 35, 36, 37, 38], they can also suffer from damping effects which leads to broadening of the spectral peaks and information loss. Here we overcome the numerical scaling issues of the KBEs by developing a new scheme that picks up features beyond the mean-field level, demonstrates excellent agreement with exact results and has effectively linear scaling in the number of time-steps.
In this letter we present a route to include dynamical correlations for the time off-diagonal Green’s function without the need to explicitly perform collision integrals. To compute the non-equilibrium spectral function one needs the off-diagonal components of the GF within a small neighborhood of the time-diagonal, centered around some time $t_{p}$, as shown schematically in Fig. 1. Here, the scaling for computing the non-equilibrium spectral function is at best $O(\tilde{N}_{t}^{2})$ , where $\tilde{N}_{t}$ is the number of time-points centered around the probe time $t_{p}$. In practice, $\tilde{N}_{t}$ is dependent on the windowing function, $\mathcal{S}(t)$ which is typically Gaussian and further discussed in the supplemental information (SI)[26]. The time-resolved spectral function is expressed as,[27]
[…]
Where $\mathcal{S}(t)$ determines the energy and temporal resolution of $\mathcal{A}(\omega,t_{p})$ [39]. In practice $\tilde{N}_{t}\ll N_{t}$ and is constant and determined solely by $\mathcal{S}(t)$ and the overall computational scaling is thus determined solely by the time-diagonal propagation.
The real-time Dyson expansion (RT-DE)
In the SI we show that[26].
$\begin{split}&G(12)=G^{\mathrm{MF}}(12)+\int\int d3d4\medspace G^{\mathrm{MF}}(13)\Sigma[G^{\mathrm{MF}}(34)]G(42)\\ &\equiv\left[i\partial_{t_{1}}-h^{\mathrm{MF}}[G^{\mathrm{MF}}(1)]\right]G(12)\\ &\hskip 71.13188pt-\int d3\medspace\Sigma[G^{\mathrm{MF}}(13)]G(32)=\delta(12)\end{split}$ (2)
The first equation, gives the Dyson equation describing a one-shot correction on-top of a reference (arbitrary mean-field Hamiltonian (MFH)) for an arbitrary self-energy, $\Sigma(t,t^{\prime})$[40]. The second is exactly the equation we will solve in this paper using a real-time GF formalism. At equilibrium, our approach reduces to the standard Many-body perturbation theory (MBPT) correction on top of a MFH. This typically leads to great improvement and even yields spectra containing signatures of strong dynamical correlations[41, 42, 43]. Equation (2) builds an intuition behind the procedure: we use a mean-field method as an initial guess and we add a correction term that includes dynamical correlation effects through $\Sigma(t,t^{\prime})$. RT-DE thus represents a natural extension of the one-shot correction in MBPT to non-equilibrium settings.
Using this form of the integro-differential equation for $G^{\mathrm{R}}(t,t^{\prime})$ and the equations of motion for $G^{\mathrm{MF}}(t,t^{\prime})$ we derive an ordinary differential equation (ODE) scheme that involves propagating $G^{\mathrm{R}}(t,t^{\prime})$ along with an auxiliary two-body propagator. A similar approach has also been applied to the time-diagonal GF in a linear scaling HF-GKBA scheme[13, 22]. Propagating the ODE scheme on the time off-diagonal for some $\tilde{N_{t}}$ steps scales as $O(\tilde{N}_{t})$. Thus reconstructing the off-diagonal from $\tilde{N}_{t}$ steps along the diagonal will indeed scale as $O(\tilde{N}^{2}_{t})$. Furthermore, the time evolution is independent for each starting point, making this scheme extremely easy to parallelize in time coordinates.
In the following, we only show the equations for $G^{\mathrm{<}}(t,t^{\prime})$ however we perform the same procedure for $G^{\mathrm{R}}(t,t^{\prime})$. The details are in the SI, which includes the derivation of the RT-DE scheme[26].
The RT-DE involves solving the following equation for $G^{<}(t,t^{\prime})$.
$\begin{split}\frac{\mathrm{d}G^{\mathrm{<}}(t,t^{\prime})}{\mathrm{d}t}&=-i\left[h^{\mathrm{MF}}(t)G^{\mathrm{<}}(t,t^{\prime})+I^{\mathrm{<}}(t,t^{\prime})\right],\\ I^{\mathrm{<}}_{im}(t,t^{\prime})&=-\sum_{klp}w_{iklp}(t)\mathcal{F}_{lpmk}(t,t^{\prime}).\\ \end{split}$ (3)
Here, the lesser collision integral, $I^{<}(t,t^{\prime})$, depends on the interaction tensor and a two-body propagator $\mathcal{F}_{ijkl}$. By employing the approximation in equation (2) we introduce an ODE for $\mathcal{F}_{ijkl}$ given by
$\begin{split}\frac{\mathrm{d}\mathcal{F}_{lpmk}(t,t^{\prime})}{\mathrm{d}t}&=\sum_{qrsj}w_{qrsj}(t)\Big{[}G^{>,\textrm{MF}}_{lq}(t)G^{<,\textrm{MF}}_{pr}(t)G^{>,\textrm{MF}}_{sk}(t)-G^{<,\textrm{MF}}_{lq}(t)G^{<,\textrm{MF}}_{pr}(t)G^{<,\textrm{MF}}_{sk}(t)\Big{]}G^{\mathrm{<}}_{jm}(t,t^{\prime})\\ &\hskip 142.26378pt-i\sum_{x}\Big{[}h^{\mathrm{MF}}_{lx}(t)\mathcal{F}_{xpmk}(t,t^{\prime})+h^{\mathrm{MF}}_{px}(t)\mathcal{F}_{lxmk}(t,t^{\prime})-\mathcal{F}_{lpmx}(t,t^{\prime})h^{\mathrm{MF}}_{xk}(t)\Big{]}.\\ \end{split}$ (4)
To fully demonstrate the the quality of this “one-shot correction” at equilibrium, we provide its numerical analysis in the SI, where we see excellent results[26]. Additionally we discuss the failings of mean-field approaches like HF and the HF-GKBA
In the non-equilibrium setting the RT-DE requires a time evolution of $G^{<}(t,t)$ using a time-dependent MFH. While our scheme is applicable to any particular choice of the reference static Hamiltonian, such as from a time-dependent density functional theory calculation[44] and even those beyond mean-field e.g from some form of a downfolding procedure, we apply TD-HF on the time-diagonals for simplicity. This is discussed further in the SI[26]. For completeness, we also show in [26] that the difference between exact diagonalization and TD-HF is relatively small on the diagonal. The time-diagonal components are inserted in equation (4). Starting from some $t^{\prime}=t$ we can time-step in the $t$ variable in the range $t^{\prime}\leq t\leq T_{\mathrm{max}}$. Here we again emphasize an important advantage of the RT-DE scheme. Since we can start from any point on the diagonal we choose to only compute those $\tilde{N}_{t}^{2}$ off-diagonal terms within the probe window.
Finally, we we need to select an initial condition for $I^{<}(t,t)$ and $G^{<}(t,t)$ which, in principle, can break the equivalence to the standard perturbative expansion. This matter is discussed in more detail in the SI, as well as the initial conditions for the retarded component[26]. For the results shown here we have taken $I^{<}(t,t)=0$ and $G^{<}(t,t)=G^{<,\mathrm{MF}}(t)$.
Application to non-equilibrium problems
We will now proceed by studying $\mathcal{A}(\omega,t_{p})$ with our proposed scheme. In this manuscript we only discuss comparisons of $\mathcal{A}(\omega,t_{p})$ since this is the experimentally relevant quantity extracted from the time off-diagonal GF. Further, this serves to compress the information stored in the GF making our comparisons clearer.
$\begin{split}\mathcal{H}&=-J\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}+U\sum_{i}n_{i\uparrow}n_{i\downarrow}+\sum_{ij}h_{ij}^{\mathrm{N.E}}(t)c_{i}^{\dagger}c_{j}\\ \end{split}$ (5)
First, to demonstrate the accuracy of our method, we apply the RT-DE to non-equilibrium system and compare to TD-HF and exact diagonalization. We excite the system given by equation (5) with a short wavelength pulse given by $h_{ij}^{\mathrm{N.E}}(t)=\delta_{ij}E\cos\left(\frac{\pi r_{i}}{2}\right)\exp\left({-\frac{(t-t_{0})^{2}}{2T_{p}^{2}}}\right)$[17], where $r_{i}=\frac{N_{s}-1}{2}-i$ is the lattice vector for site $i$ and $N_{s}$ is the number of lattice sites. Results for the lesser and retarded spectrum for dynamics induced by a long wavelength electric field and by a quench of the system are discussed in the SI[26].
Figure 2: Time-resolved emission spectrum exact diagonalization (ED), RT-DE which is in excellent agreement with ED, and HF-GKBA that mostly copies the TD Hartree Fock (TD-HF) spectrum. Calculated in the model given in equation (5) with the following parameters: $U=1.0J$, $N_{s}=4$, $E=0.5J$, $t_{0}=5J^{-1}$ and $T_{p}=0.5J^{-1}$ after time evolving to $T_{\mathrm{max}}=200J^{-1}$. The probe width is taken to be $\delta=10$ and $t_{p}=100J^{-1}$.
Examining Fig. 2 we see the RT-DE is in excellent agreement with the exact result for the lesser component of the $\mathcal{A}(\omega,t_{p})$. It picks up features around $\omega\approx-1.2J$ and $\omega\approx-0.4J$ that are completely missed by TD-HF and HF-GKBA, which mostly copies the TD-HF features. Notably, neither of these peaks is present in equilibrium[26]. This shows the RT-DE can pick up emergent features that arise purely due to non-equilibrium effects in the system. For the peaks above the Fermi energy,the RT-DE, HF-GKBA and TD-HF pick up the peaks at $\omega\approx 1.0J$ and $2.0J$. TD-HF and the RT-DE methods slightly underestimate the energy of the peak around $\omega\approx 1.0J$ by the same amount. Interestingly, the peak near $\omega\approx 2.0J$ is capture slightly better with TD-HF than with the RT-DE. This is related to the initial condition for the lesser component. Given that the initial condition for $G^{<}(t,t^{\prime})$ can be fixed in different ways. Further optimization and constraining the particular choices will be explored elsewhere. However, this boundary condition affects only the details of the electronic distribution in dynamically renormalized states. The band renormalization and emergence of new spectral peaks is however captured perfectly using the knowledge of the $G^{R}$ alone, which is free of any uncertainty. Finally, we note the appearance of two additional peaks in the HF-GKBA result at $\omega\approx 0.5J$ and $\omega\approx 1.6J$
For completeness we provide discussion of the full spectrum from $G^{\mathrm{R}}(t,t^{\prime})$ in the SI[26].
In our second example we explore a realistic model of an insulator out of equilibrium. We study the dynamics of a long-range interacting two band Hubbard model with periodic boundary conditions at half filling under a strong optical excitation. The full Hamiltonian is given by
$\begin{split}\mathcal{H}&=\sum_{\alpha,\beta\in\{c,v\}}\sum_{\langle i,j\rangle,\sigma}h^{\alpha\beta}_{ij}(t)c^{\dagger\alpha}_{i,\sigma}c^{\beta}_{j,\sigma}+U\sum_{i,\alpha}n^{\alpha}_{i\uparrow}n^{\alpha}_{i\downarrow}\\ &\hskip 71.13188pt+U\sum_{i}n_{ic}n_{iv}+\gamma U\sum_{i<j\atop\alpha\leq\beta}\frac{n_{i\alpha}n_{j\beta}}{|\vec{r}_{i}-\vec{r}_{j}|}.\\ h^{\alpha\beta}_{ij}(t)&=J\delta_{\alpha\beta}\delta_{\langle i,j\rangle}(\delta_{\alpha v}-\delta_{\alpha c})+\epsilon_{\alpha}\delta_{\alpha\beta}\\ &\hskip 42.67912pt+\delta_{ij}(1-\delta_{\alpha\beta})E\cos(\omega_{p}(t-t_{0}))\mathrm{e}^{-\frac{(t-t_{0})^{2}}{2T_{p}^{2}}}\end{split}$ (6)
Figure 3: Non-equilibrium time-resolved emission spectrum for the model given in equation (6) with 12 sites evolved until time $T_{\mathrm{max}}=100J^{-1}$. Panels a), b) and c) show the spectra computed with TD-HF, the HF-GKBA and the RT-DE respectively. The model parameters are $U=4.0J$, $\gamma=0.5$, $\epsilon_{v}=-30J$, $\epsilon_{c}=-20J$, $E=2.5J$, $\omega_{p}=7.5J$, $t_{0}=10J^{-1}$, $T_{p}=1.8J^{-1}$. The probe width used in (1) is $\delta=5$ and the probe is centered at $t_{p}=90J^{-1}$. A prominent excitonic feature is seen for the RT-DE results just below the renormalized conduction band; further, a visible band satellite is visible at the bottom of the bandstructure.
The results are shown in Fig. 3. We compare TD-HF, the HF-GKBA and the RT-DE using the following model parameters: $U=4.0J$, $\epsilon_{v}=-30J$, $\epsilon_{c}=-20J$, $\gamma=0.5,E=2.5J$, $\omega_{p}=7.5J$, $t_{0}=10J^{-1}$ and $T_{p}=1.8J^{-1}$. The excitation frequency (pump), $\omega_{p}$ is below the equilibrium gap ($\sim 10.0J$) but it is chosen such that excitons can form in the system. We also remark that at equilibrium the three methods produce practically identical band structures, shown as red dashed lines in Fig. 3. This allows us to see how each method performs, purely from changes due to the system leaving equilibrium.
The periodic system is too large for benchmarking against exact results; hence, this example serves primarily as a demonstration of the ability of the RT-DE to be applied to relatively large periodic systems. Even in the absence of the exact reference, we compare the qualitative features of $\mathcal{A}(\omega,t_{p})$ and determine the physical origins of the dynamical features that are produced by RT-DE (but not by the other methods).
Fig. 3 a) shows $\mathcal{A}(\omega,t_{p})$ for TD-HF. The full non-equilibrium band structure is shown in blue dash-dotted lines. Due to the lack of dynamical self-energy (yielding possibly multiple quasiparticle solutions) no new (satellite) states can appear in the TD-HF result. The figure shows electrons are excited across the gap and into the conduction band around $k=0$. This change in the electron distribution leads to a renormalization of the band gap, which is governed by the changes in the orbital occupancies and related to changes in the time dependent density matrix. In practice, occupied states are shifted down in energy creating a relatively sharp discontinuity in the upper band which becomes partially filled around $k=0$. The concentration of electrons (holes) at the the conduction (valence) band extrema and the mean-field nature of TD-HF means the renormalization mainly effects the $k=0$ point, leading to a significant stretching of the bands. As expected the lack of two-body correlations in TD-HF means we observe no novel features in $\mathcal{A}(\omega,t_{p})$ which could be interpreted as emerging excitonic states beyond the band stretching.
In Fig. 3 b) we again show $\mathcal{A}(\omega,t_{p})$, now computed using the HF-GKBA. This result shares features common to TD-HF as we expect because it is dominated by the off-diagonal terms dominated by HF. Firstly, while the HF signatures are diminished, we see a similar stretching of the bands relative to their equilibrium value. One reason for this change comes from an inconsistency between the treatment of the time-diagonal and off-diagonal terms in HF-GKBA, as the HF-GKBA density matrix, $\rho^{\mathrm{HF-GKBA}}(t)$, enters the HF equation for the off-diagonal instead of $\rho^{\mathrm{HF}}(t)$. The density matrix is thus projected on to a set of excited states of the HF Hamiltonian. Such an inconsistency makes it a priori difficult to interpret HF-GKBA spectral results. Importantly the HF-GKBA again lacks additional excitonic features. The picture is thus consistent with TD-HF (albeit with distinct band renormalization ).
Finally we turn to Fig. 3 c), showing the RT-DE results. The results differ markedly and show emergent features. For the main renormalized bands we see a renormalization of the band positions compared to equilibrium. This affects both the $k\neq 0$ and $k=0$ points very similarly resulting in almost a rigid shift of the bands, compared to the stretching seen in Fig 3 a) and b). This is consistent with renormalized (increased) screening in the photoexcited system affecting the entire band structure and not only the small number states that become occupied by photo-excitation.
In addition, there are two more striking differences. The first is a peak appearing just below the renormalized conduction band, consistent with an in-gap electronic state, i.e., an exciton. Secondly we see the appearance of a satellite band appearing just below the position of the equilibrium valence band. Similarly to the finite system in Fig. 2 the inclusion of dynamical correlations through $I^{<}(t,t^{\prime})$ allows for the formation of additional spectral features. This is consistent with satellites emerging due to coupling to a bosonic excitation in the system. Here, the satellite is displaced by energy equivalent to the distance between the excitonic feature and the minimum of the renormalized conduction state. Potentially, one can interpret this as an “excitonic replica”, i.e., simulaneous QP formation and exciton dissociation, similar to other bosonic replicas[42]. The exact interpretation of such features are left for future research, but we note that these are consistent with the exact solutions made for finite systems studied above.
Conclusions and outlook
For our tests in finite systems the results for the RT-DE in and out of equilibrium are in excellent agreement with exact results. Not only does the RT-DE correct features picked up by TD-HF and the HF-GKBA but it also captures satellite features that both other methods miss entirely. It is important to highlight that these excellent results are produced using the second Born approximation (representing one of the simplest self-energy forms). Thus, the results and comparisons to exact diagonalization given in this manuscript serve as compelling a proof of principle of the RT-DE. Already our method can capture dynamical correlation effects such as plasmons to a high degree of accuracy and it is natural to expect that extending to more advanced self-energy approximations will only further improve upon the results presented here, an area that is currently being pursued. To show the generality of our approach we have provided a derivation of the RT-DE for the $GW$ self-energy in the SI[26], it’s implementation, testing and comparison is left to a later publication.
For our periodic system, the RT-DE provides physically meaningful results. Importantly by including two-body correlations through $\Sigma(t,t^{\prime})$ we are able to observe an emergent excitonic peak appearing in $\mathcal{A}(\omega,t_{p})$ and energetically well separated from the quasiparticle band.
In conclusion the RT-DE is an efficient scheme for studying the non-equilibrium spectral properties of quantum systems, offering accuracy and a huge reduction in cost relative to the full KBEs. While KBEs and RT-DE will scale similarly in system size, by reducing the cost from cubic to linear in the number of time-steps, we overcome one of the major limitations of the KBEs that has prevented them being widely applied to non-equilibrium first principles simulations. Application of this scheme to a more realistic system and comparison of the results to experimental measurements as a validation is underway. Finally, we note that in principle the RT-DE can be combined with existing real-time MBPT methods and their extension to more advanced self-energy approximations, such as $GW\Gamma$[41]. This extension is currently underway and will open the doors to the accurate and efficient simulation of time-resolved photo-emission experiments.
Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program under Award Number DE-SC0022198. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award BES-ERCAP0029462